# A torsion theoretic interpretation for sheaves of modules and Grothendieck topologies on directed categories

### v2

@@ 7. Applications
In this section we describe some potential applications of our results in representation theory of combinatorial categories and groups.
7.1. Local self-injectivity of infinite full subcategories
Let ${\mathscr{C}}$ be the category of finite sets and injections, or the category of finite dimensional vector spaces over a finite field $\mathbb{F}_{q}$ and $\mathbb{F}_{q}$-linear injections. It has been proved in [11] (see also [22]) that every indecomposable projective ${\mathscr{C}}$-module over a field $k$ of characteristic 0 is also injective. In this subsection we extend this result to their infinite full subcategories.
Recall that an EI category ${\mathscr{C}}$ of type $\mathbb{N}$ is called locally finite if ${\mathscr{C}}(m,n)$ is a finite set for $m,n\in\mathbb{N}$.
Theorem 7.1.
Let ${\mathscr{C}}$ be a locally finite EI category of type $\mathbb{N}$ such that every morphism is monic, ${\mathscr{D}}$ an infinite full subcategory of ${\mathscr{C}}$, and $k$ a field. If every indecomposable projective ${\mathscr{C}}$-module over $k$ is injective, then so is every indecomposable projective ${\mathscr{D}}$-module over $k$.
@@ 7.1. Local self-injectivity of infinite full subcategories
Since ${\mathscr{D}}$ is a full subcategory of ${\mathscr{C}}$, one may impose a rigid Grothendieck topology $J$ on ${\mathscr{C}}^{{\mathrm{op}}}$ such that $J$-irreducible objects in ${\mathscr{C}}$ are precisely objects in ${\mathscr{D}}$. Thus by Remark 5.17, one has
@@
${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J,\,\underline{k})\simeq{\mathscr{D}}\operatorname{-Mod},$
where $\underline{k}$ is the constant structure sheaf on the site $({\mathscr{C}}^{{\mathrm{op}}},J)$, and ${\mathscr{D}}\operatorname{-Mod}$ is the category of ${\mathscr{D}}$-modules over $k$.
Since ${\mathscr{D}}$ is locally finite, to establish the conclusion, it suffices to show that $k{\mathscr{D}}(n,-)$ is injective for each object $n$ in ${\mathscr{D}}$. Note that by [13, C2.2, Theorem 2.2.3], the equivalence from the sheaf category to the module category is given by the restriction functor; and furthermore the restriction of $P(n)=k{\mathscr{C}}(n,-)$ is precisely $k{\mathscr{D}}(n,-)$. Consequently, one only needs to show that $P(n)$ is an injective object in ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J,\,\underline{k})$. By Corollary 4.4, this is equivalent to saying that $P(n)$ is a $J$-torsion free injective ${\mathscr{C}}$-module. But we have assumed that $P(n)$ is an injective ${\mathscr{C}}$-module, so it remains to show that $P(n)$ is $J$-torsion free.
Since ${\mathscr{C}}$ is an infinite category, $J$ must be a generic Grothendieck topology by Proposition 6.6. Thus $J(n)\subseteq J_{a}(n)$ for $n\in\mathbb{N}$, where $J_{a}$ is the atomic Grothendieck topology. Accordingly, one has $\mathcal{T}(J)\subseteq\mathcal{T}(J_{a})$ by Lemma 3.6, so every $J_{a}$-torsion free ${\mathscr{C}}$-module over $k$ is also $J$-torsion free. Clearly, $P(n)$ is $J_{a}$-torsion free as every morphism in ${\mathscr{C}}$ is monic, so it is $J$-torsion free as well. This finishes the proof. ∎
@@ 7.2. Representations of groups
Let $G$ be a group, and $\mathcal{H}$ a family of subgroups of $G$. The orbit category ${\mathrm{Orb}}(G,\mathcal{H})$ of $G$ with respect to $\mathcal{H}$ is defined as follows: objects are left cosets $G/H$ with $H\in\mathcal{H}$, and morphisms from $G/H$ to $G/K$ are $G$-equivariant maps. Every such morphism is an epimorphism, and can be represented by an element $g\in G$ such that $gHg^{-1}\subseteq K$. In particular, for every object $G/H$, its endomorphism monoid is a group isomorphic to $N_{G}(H)/H$, so ${\mathrm{Orb}}(G,\mathcal{H})$ is an EI category. For more details, the reader can refer to [18, III.9]. Moreover, it is easy to see that ${\mathrm{Orb}}(G,\mathcal{H})$ is noetherian (resp., artinian) if and only if the poset $\mathcal{H}$ with respect to inclusion is noetherian (resp., artinian).
@@
Let $G$ be a topological group, and let $\mathcal{H}$ be a fundamental system of open subgroups. Impose the atomic Grothendieck topology $J_{a}$ on ${\mathrm{Orb}}(G,\mathcal{H})$, and let $\underline{k}$ be a constant structure sheaf given by a commutative ring $k$. Artin’s theorem [18, III.9 Theorem 1] asserts that
@@
a result described in [26, Subsection 5.3].
@@
Let $G$ be a finite group, $p$ a prime number, and $k$ an algebraically closed field of characteristic $p$. A weight is a pair $(H,L)$ with $H$ a $p$-subgroup (possibly trivial) of $G$ and $L$ is an irreducible projective $kN_{G}(H)/H$-module. Two weights $(H_{1},L_{1})$ and $(H_{2},L_{2})$ are equivalent if there exists $g\in G$ such that the conjugate $H_{1}^{g}$ coincides with $H_{2}$ and $L_{1}^{g}\cong L_{2}$. Alperin’s weight conjecture asserts that the number of isomorphism classes of irreducible $kG$-modules is equal to the number of equivalence classes of weights.
@@
    such that $V(H,L)\cong V(H^{\prime},L^{\prime})$ if and only if $(H,L)$ and $(H^{\prime},L^{\prime})$ are equivalent?

### v3

@@ 7. Applications
In this section we describe some applications of our results in representation theory of combinatorial categories and groups.
7.1. Representations of infinite full subcategories of ${\mathrm{FI}}$ and ${\mathrm{VI}}_{q}$
Throughout this subsection let $k$ be a commutative noetherian ring, let ${\mathscr{C}}$ be a skeleton of the category ${\mathrm{FI}}$ of finite sets and injections, or the category ${\mathrm{VI}}_{q}$ of finite dimensional vector spaces over a finite field $\mathbb{F}_{q}$ and $\mathbb{F}_{q}$-linear injections, and let ${\mathscr{D}}$ be a full subcategory of ${\mathscr{C}}$ with infinitely many objects. Denote objects in ${\mathscr{C}}$ by $\boldsymbol{n}$, $n\in\mathbb{N}$, which is isomorphic to $[n]=\{1,2,\ldots,n\}$ for ${\mathrm{FI}}$ or $\mathbb{F}_{q}^{n}$ for ${\mathrm{VI}}_{q}$. Accordingly, let $P(n)$ be $k{\mathscr{C}}(\boldsymbol{n},-)$.
One may impose a rigid Grothendieck topology $J_{r}$ on ${\mathscr{C}}^{{\mathrm{op}}}$ such that $J_{r}$-irreducible objects in ${\mathscr{C}}$ are precisely objects in ${\mathscr{D}}$. On the other hand, it is easy to check that ${\mathscr{C}}^{{\mathrm{op}}}$ satisfies the right Ore condition, so one can also impose the atomic Grothendieck topology $J_{a}$ on it.
Lemma 7.1.
Notation as above. For each $n\in\mathbb{N}$, $P(n)$ is a sheaf of modules on the ringed site $({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$, where $\underline{k}$ is the constant structure sheaf, and hence a sheaf of modules on the ringed site $({\mathscr{C}}^{{\mathrm{op}}},\,J_{d},\,\underline{k})$.
@@ 7.1. Representations of infinite full subcategories of ${\mathrm{FI}}$ and ${\mathrm{VI}}_{q}$
Since ${\mathscr{D}}$ has infinitely many objects, it follows that for each object $\boldsymbol{n}$ in ${\mathscr{C}}$, there are morphisms starting from it and ending at a certain object in ${\mathscr{D}}$. Consequently, $J_{r}(n)$ is nonempty, and hence $J_{r}(n)\subseteq J_{a}(n)$ for each $n\in\mathbb{N}$. Consequently, every sheaf of modules over $({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$ is also a sheaf of modules over $({\mathscr{C}}^{{\mathrm{op}}},\,J_{d},\,\underline{k})$, so one only needs to show the first statement.
This is clear. Indeed, Artin’s theorem [23, III. 9, Theorems 1 and 2] provides an equivalence
${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})\simeq kG\operatorname{-dMod},$
[…]
Remark 7.2.
One can also prove the conclusion by showing that $P(n)$ is $J_{a}$-saturated. For ${\mathrm{FI}}$, this is established in [21, Theorem C]; for ${\mathrm{VI}}_{q}$, under the extra assumption that $q$ is invertible in $k$, this follows from [25, Corollary 4.22].
It follows from this lemma that both $J_{d}$ and $J_{a}$ are subcanonical; that is, representable presheaves are sheaves. Therefore, the following set is a family of generators of ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$:
$\{P(n)\mid n\in\mathbb{N}\},$
so we can make the following definition.
Definition 7.3.
We say that $V\in{\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$ is finitely generated if there is a surjection
$\bigoplus_{n}P(n)^{c_{n}}\twoheadrightarrow V$
such that the sum of all multiplicities $c_{n}$ is finite. Finitely generated sheaves in ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{d},\,\underline{k})$ are defined similarly.
In the rest of this subsection we will frequently use the following easy observation: a sheaf homomorphism $V\to W$ is injective (resp., surjective) if and only if it is injective (resp., surjective) as a presheaf homomorphism. Consequently, $V\in{\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$ is finitely generated if and only if viewed as a ${\mathscr{C}}$-module it is finitely generated. The same conclusion holds for ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{d},\,\underline{k})$.
Proposition 7.4.
Every finitely generated ${\mathscr{D}}$-module over $k$ is noetherian.
Proof.
It suffices to prove the following conclusion: for each $\boldsymbol{n}\in\operatorname{Ob}({\mathscr{D}})$, every submodule $V$ of $Q(n)=k{\mathscr{D}}(\boldsymbol{n},-)$ is finitely generated. By [17, C2.2, Theorem 2.2.3] and Remark 5.19, one has
${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{r},\,\underline{k})\simeq{\mathscr{D}}\operatorname{-Mod},$
for which an explicit equivalence is given by the restriction functor $\operatorname{res}$ from the sheaf category to the module category and the coinduction functor $\mathrm{coind}$ (the right Kan extension) in the other direction. Accordingly, $\operatorname{coind}(V)$ is a subsheaf of $\operatorname{coind}(Q(n))$ in ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{r},\,\underline{k})$. But $P(n)\in{\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{r},\,\underline{k})$ by Lemma 7.1 and $\operatorname{res}(P(n))=Q(n)$, it follows that $\operatorname{coind}(Q(n))\cong P(n)$. Thus without loss of generality we can assume that $\operatorname{coind}(V)$ is a subsheaf of $P(n)$.
By the noetherianity results in [7, 28], $\operatorname{coind}(V)$ is a finitely generated ${\mathscr{C}}$-module, so one has a surjection
$\bigoplus_{i\in\mathbb{N}}P(i)^{c_{i}}\longrightarrow\operatorname{coind}(V)$
of ${\mathscr{C}}$-modules such that the sum of these $c_{i}$’s is finite. By the above observation, this is also a surjection in ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{r},\,\underline{k})$. Applying the restriction functor, we obtain a surjection
$\bigoplus_{i\in\mathbb{N}}\operatorname{res}(P(i))^{c_{i}}\longrightarrow\operatorname{res}(\operatorname{coind}(V))\cong V$
of ${\mathscr{D}}$-modules. It remains to show that each $\operatorname{res}(P(i))$ is a finitely generated ${\mathscr{D}}$-module.
In the case $\boldsymbol{i}\in\operatorname{Ob}({\mathscr{D}})$, one has $\operatorname{res}(P(i))=Q(i)$, so the conclusion holds trivially. In the case $\boldsymbol{i}\notin\operatorname{Ob}({\mathscr{D}})$, let $\boldsymbol{j}$ be the minimal number such that $j>i$ and $\boldsymbol{j}\in\operatorname{Ob}({\mathscr{D}})$, which always exist since ${\mathscr{D}}$ is an infinite subcategory of ${\mathscr{C}}$. Then the ${\mathscr{D}}$-module $\operatorname{res}(P(i))$ is generated by its value on $\boldsymbol{j}$, and hence is finitely generated as well. The proof is completed. ∎
Let ${\mathrm{sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$ be the category of finitely generated objects in ${\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$. The above proposition tells us that ${\mathrm{sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$ is an abelian category; for details, see the comment before [8, Remark 3.15].
In the rest of this subsection let $k$ be a field of characteristic 0.
Proposition 7.5.
Every indecomposable projective ${\mathscr{D}}$-module over $k$ is also injective.
Proof.
It suffices to show that $Q(n)=k{\mathscr{D}}(\boldsymbol{n},-)$ is injective for each $\boldsymbol{n}\in\operatorname{Ob}({\mathscr{D}})$. Using the equivalence from ${\mathscr{D}}\operatorname{-Mod}$ to ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{r},\,\underline{k})$ given by the coinduction functor, one only needs to show the injectivity of $P(n)\cong\operatorname{coind}(Q(n))$ in ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{r},\,\underline{k})$. By Corollary 4.4, this is equivalent to saying that $P(n)$ is a $J_{r}$-torsion free injective ${\mathscr{C}}$-module. But it has been proved in [12] that $P(n)$ is an injective ${\mathscr{C}}$-module, so it remains to show that $P(n)$ is $J_{r}$-torsion free. This is clear. Indeed, $P(n)$ is $J_{a}$-torsion free as every morphism in ${\mathscr{C}}$ is monic, so it is $J_{r}$-torsion free as well by Lemma 3.6. This finishes the proof. ∎
To prove other statements in Theorem 1.8, we turn to the atomic Grothendieck topology $J_{a}$ on ${\mathscr{C}}^{{\mathrm{op}}}$. Since ${\mathscr{D}}$ has infinitely many objects, ${\mathscr{D}}^{{\mathrm{op}}}$ is a $J_{a}$-dense subcategory in the sense of [17, C2.2, Definition 2.2.1]. Furthermore, the induced Grothendieck topology $K$ on ${\mathscr{D}}^{{\mathrm{op}}}$ via setting
$K(n)=\{S\cap{\mathscr{D}}(\boldsymbol{n},-)\mid S\in J_{a}(n)\}$
coincides with the atomic Grothendieck topology. Thus by abuse of notation we also use $J_{a}$ to denote this induced Grothendieck topology. Applying [17, C2.2, Theorem 2.2.3], we obtain an immediate result:
Lemma 7.6.
There is an equivalence
${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})\simeq{\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k}).$
given by the restriction functor and the coinduction functor.
Remark 7.7.
[…]
$\{H_{n}=\{g\in G\mid g(i)=i,\,\forall i\in[n]\}\mid\boldsymbol{n}\in\operatorname{Ob}({\mathscr{C}})\}$
and
$\{H_{n}=\{g\in G\mid g(i)=i,\,\forall i\in[n]\}\mid\boldsymbol{n}\in\operatorname{Ob}({\mathscr{D}})\}$
are cofinal systems of open subgroups of $G$, one has
${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})\simeq kG\operatorname{-dMod}\simeq{\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k}).$
For details, please refer to [8, Section 5].
The above equivalence gives another proof of Proposition 7.5. For $\boldsymbol{n}\in\operatorname{Ob}({\mathscr{D}})$, since $P(n)$ is a $J_{a}$-torsion free injective ${\mathscr{C}}$-module, it is also injective in ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$ by Corollary 4.4. Applying the restriction functor, we deduce that $Q(n)\cong\operatorname{res}(P(n))$ is injective in ${\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$. Again, by Corollary 4.4, $Q(n)$ is an injective ${\mathscr{D}}$-module.
Lemma 7.8.
Let $V$ be a finitely generated $J_{a}$-torsion free ${\mathscr{D}}$-module. Then there is an injection $V\to\operatorname{res}(P)$ with $P$ a finitely generated projective ${\mathscr{C}}$-module.
Proof.
Since $V$ is $J_{a}$-torsion free, one may obtain an injection $V\to V^{\sharp}$ in ${\mathscr{D}}\operatorname{-Mod}$, where $V^{\sharp}$ is the sheafification of $V$ with respect to $J_{a}$. Thus without loss of generality we can assume that $V$ is contained in ${\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$. Accordingly, $\operatorname{coind}(V)$ is contained in ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$.
By [12, Proposition 7.5] and [25, Theorem 4.34], there is an injection $\operatorname{coind}(V)\to P$ in ${\mathscr{C}}\operatorname{-Mod}$, where $P$ is a finitely generated projective ${\mathscr{C}}$-module. This is also an injection in ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$ by the above mentioned observation. By Lemma 7.6, one obtains an injection $V\to\operatorname{res}(P)$ in ${\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$, which is the desired injection in ${\mathscr{D}}\operatorname{-Mod}$. ∎
Now we finish the proof of Theorem 1.8.
Proposition 7.9.
Let $k$ be a field of characteristic 0. Then:
  1. (1)
    the category ${\mathscr{D}}\operatorname{-mod}$ of finitely generated ${\mathscr{D}}$-modules has enough injectives;
  2. (2)
    up to isomorphism, an indecomposable injective ${\mathscr{D}}$-module is one of the following three:
    • a summand of $k{\mathscr{D}}(\boldsymbol{n},-)$ for a certain $\boldsymbol{n}\in\operatorname{Ob}({\mathscr{D}})$,
    • a summand of the restriction of $k{\mathscr{C}}(\boldsymbol{n},-)$ for a certain $\boldsymbol{n}\in\operatorname{Ob}({\mathscr{C}})\setminus\operatorname{Ob}({\mathscr{D}})$,
    • a finite dimensional injective ${\mathscr{D}}$-module;
  3. (3)
    one has the following equivalences
    ${\mathscr{D}}\operatorname{-mod}/{\mathscr{D}}\operatorname{-fdmod}\simeq{\mathscr{C}}\operatorname{-fdmod}.$
  4. (4)
    every finitely generated ${\mathscr{D}}$-module has finite injective dimension.
Proof.
(1). A finitely generated ${\mathscr{D}}$-module $V$ gives a short exact sequence
$0\to V_{T}\to V\to V_{F}\to 0$
in ${\mathscr{D}}\operatorname{-Mod}$, where $V_{T}$ is $J_{a}$-torsion and $V_{F}$ is $J_{a}$-torsion free. Since $V_{T}$ is finitely generated, it is finite dimensional. Hence, $V_{T}$ can be embedded into a finite dimensional injective ${\mathscr{D}}$-module. By Lemma 7.8, $V_{F}$ can be embedded into $\operatorname{res}(P)$ with $P$ a finitely generated projective ${\mathscr{C}}$-module. But $P$ is injective in ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$, so $\operatorname{res}(P)$ is injective in ${\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$, and hence injective in ${\mathscr{D}}\operatorname{-Mod}$.
(2). Let $V$ be an indecomposable injective ${\mathscr{D}}$-module. By (1), it can be embedded into a direct sum of a finite dimensional injective ${\mathscr{D}}$-module and $\operatorname{res}(P)$, so it is either finite dimensional or a direct summand of $\operatorname{res}(P)$. In the later case, it must be a direct summand of some $\operatorname{res}(P(n))$. The conclusion then follows.
(3). By [8, Corollary 4.9], one has
${\mathrm{sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})\simeq{\mathscr{C}}\operatorname{-mod}/{\mathscr{C}}\operatorname{-fdmod}\simeq{\mathscr{C}}\operatorname{-fdmod}$
The finitely generated version of [8, Corollary 3.9] gives another equivalence
${\mathrm{sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})\simeq{\mathscr{D}}\operatorname{-mod}/{\mathscr{D}}\operatorname{-fdmod}.$
But it is clear that the equivalence in Lemma 7.6 restricts to an equivalence
${\mathrm{sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})\simeq{\mathrm{sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k}).$
The conclusion then follows.
(4). We can use the same technique as that of the proof of [27, Thorem 4.3.1]. Let $V$ be a finitely generated ${\mathscr{D}}$-module. Since ${\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})\simeq{\mathscr{D}}\operatorname{-mod}/{\mathscr{D}}\operatorname{-fdmod}$, the kernel and cokernel of the natural map from $V$ to its sheafification $V^{\sharp}$ in ${\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$ are of finite dimensional. But every finite dimensional ${\mathscr{D}}$-module has finite injective dimension. Therefore, we only need to show that $V^{\sharp}$ has finite injective dimension in ${\mathscr{D}}\operatorname{-mod}$.
@@
Since ${\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})\simeq{\mathscr{C}}\operatorname{-fdmod}$, $V^{\sharp}$ has a finite injective resolution
$V^{\sharp}\to I^{0}\to\ldots\to I^{s}\to 0$
in ${\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,J_{a},\,\underline{k})$. This is a finite complex in ${\mathscr{D}}\operatorname{-mod}$ for which all terms except $V^{\sharp}$ are injective. Moreover, all cohomology groups are finite dimensional ${\mathscr{D}}$-modules. Consequently, $V^{\sharp}$ as a ${\mathscr{D}}$-module has finite injective dimension as desired. ∎
Relations among these categories are illustrated in the following commutative diagram
[…]
where $\mathrm{loc}$ is the localization functor, and $\bar{\nu}$ is the functor induced by the Nakayama functor $\nu$ (see [14] for details).
Remark 7.10.
It is well known that every finitely generated $J_{a}$-torsion free injective ${\mathscr{C}}$-module is projective. This is not the case in ${\mathscr{D}}\operatorname{-Mod}$. Moreover, the functor
$\bar{\nu}:{\mathscr{D}}\operatorname{-mod}/{\mathscr{D}}\operatorname{-fdmod}\longrightarrow{\mathscr{D}}\operatorname{-fdmod}$
is not an equivalence. For example, let ${\mathscr{D}}$ be obtained via removing the object $\boldsymbol{0}$ from ${\mathscr{C}}$. Then the constant ${\mathscr{D}}$-module $\underline{k}$ is injective and $J_{a}$-torsion free, but not projective. Furthermore, it was sent to the zero module in ${\mathscr{D}}\operatorname{-fdmod}$ via $\bar{\nu}$.
Remark 7.11.
Based on the work of Güntürkün-Snowden [16] and [8, 13], one can use sheaf theory to deduce similar results for the category ${\mathrm{OI}}$ of finite linearly ordered sets and order-preserving injections. For instances, if $k$ is a field of characteristic 0, and ${\mathscr{D}}$ is an infinite full subcategory of a skeleton ${\mathscr{C}}$ of ${\mathrm{OI}}$, then one has
${\mathscr{D}}\operatorname{-mod}/{\mathscr{D}}\operatorname{-tmod}\simeq{\mathscr{C}}\operatorname{-fdmod},$
where ${\mathscr{D}}\operatorname{-tmod}$ is the category of finitely generated $J_{a}$-torsion ${\mathscr{D}}$-modules. Note that in this case finite dimensional modules are $J_{a}$-torsion, but not all finitely generated $J_{a}$-torsion modules are finite dimensional.
@@ 7.2. Representations of groups
Let $G$ be a group, and $\mathcal{H}$ a family of subgroups of $G$. The orbit category ${\mathrm{Orb}}(G,\mathcal{H})$ of $G$ with respect to $\mathcal{H}$ is defined as follows: objects are left cosets $G/H$ with $H\in\mathcal{H}$, and morphisms from $G/H$ to $G/K$ are $G$-equivariant maps. Every such morphism is an epimorphism, and can be represented by an element $g\in G$ such that $gHg^{-1}\subseteq K$. In particular, for every object $G/H$, its endomorphism monoid is a group isomorphic to $N_{G}(H)/H$, so ${\mathrm{Orb}}(G,\mathcal{H})$ is an EI category. For more details, the reader can refer to [23, III.9]. Moreover, it is easy to see that ${\mathrm{Orb}}(G,\mathcal{H})$ is noetherian (resp., artinian) if and only if the poset $\mathcal{H}$ with respect to inclusion is noetherian (resp., artinian).
@@
Let $G$ be a topological group, and let $\mathcal{H}$ be a cofinal system of open subgroups. Impose the atomic Grothendieck topology $J_{a}$ on ${\mathrm{Orb}}(G,\mathcal{H})$, and let $\underline{k}$ be a constant structure sheaf given by a commutative ring $k$. Artin’s theorem [23, III.9 Theorem 1] asserts that
@@
a result described in [32, Subsection 5.3].
@@
[…]
@@
[…]